OpenStudy (anonymous):

[QUICK ANSWERS PLEASE] In the equation p=-40q+163, which relationship between p and q is true? A. You can't determine the relationship based on this info. B. q is dependent on p. C. p & q are independent of each other. D. p is dependent on q. PLEASE QUICK ANSWERS

OpenStudy (anonymous):

D

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

Yes, because if you change the q value, your p value will change. p=-40(0)+163 p=163 p=-40(1)+163 p=123

OpenStudy (anonymous):

thanks :)